OpenStudy (anonymous):

how can i integrate sin x/((cos x)^5)

OpenStudy (anonymous):

Is this the function? \[f(x)=\frac{ \sin(x) }{ \cos^{5}(x) }\]

OpenStudy (anonymous):

sin(x)/cos(x)=tan(x) 1/cos(x)^4=sec(x)^4 u=sec(x) derivative of sec(x)=sec(x) * tan(x) du=sec(x) * tan(x) int(u^3*du) (u^4)/4 (sec(x)^4)/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool, I wrote it out but got beaten to the punch. But that answer is correct, make sense?

OpenStudy (anonymous):

Also, you could have made u=cos(x) to save you a step.

OpenStudy (anonymous):

ok john, is notme's step correct?

OpenStudy (anonymous):

It is correct, but most people don't know the derivative of sec(x) off-hand so I found it easier to work this way. Both will yield the same solution, but it does show the importance of knowing your trig derivatives/integrals and how they come to be. Here's my solution for reference.

OpenStudy (anonymous):

thank you bro

OpenStudy (anonymous):

Glad to help